OpenStudy (anonymous):

What is the equation of the line that passes through ( 1,5) and (4,11)? Select one: a. y = 6 b. y = x + 2 c. y = 2x + 3 d. y = 3x + 4 ANSWER PLEASE

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

To find the slope: (y-y)/(x-x) =(11-5)/(4-1) =6/3 =2 So you know so far that the equation is y=2x+__ Just plug in the x and y value of any of the points into the partially completed equation. (5)=2(1)+3 or (11)=2(4)+3 The equation is y=2x+3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):